OpenStudy (anonymous):

Find the inverse function for f(t) = 88e^0.25t

OpenStudy (zzr0ck3r):

y = 88e^((1/4)t) we need to solve for t devide away the 88 y/88 = e^((1/4)*t) ln(y/88) = (1/4)t t = 4ln(y/88) so the inverse is y = 4ln(t/88)

OpenStudy (anonymous):

thanks